OpenStudy (anonymous):

What is the radius of a circle x^2 + y^2 +8x -10y -9 =0

OpenStudy (anonymous):

compare with \[x^2+y^2+2gx+2fy+c=0\] 2g=8,g=4 2f=-10 f=-10/2=-5 centre is(-g,-f) i.e.,(-4,5) or complete the squares and compare with \[(x-h)^2+(y-k)^2=r^2,centre~ is ~ (h,k)\]

OpenStudy (anonymous):

Would the correct answer be 5 [\sqrt{2}\]@surjithayer

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

by one method centre is (-4,5) and radius=\[\sqrt{g^2+f^2-c}=\sqrt{\left( -4 \right)^2+\left( 5 \right)^2-(-9)}=\sqrt{16+25+9}=\sqrt{50}=5\sqrt{2}\] you are correct.